# Relaying Quantum Information

### v1

@@ Abstract
Quantum relays are central to both quantum communication and distributed quantum computing, enabling long-distance transmission and modular architectures. Unlike classical repeaters, quantum repeaters preserve coherence without amplifying quantum information, relying on entanglement swapping and quantum error correction to overcome loss and decoherence. In this work, we investigate the transmission of quantum information via quantum relay channels. Our three-terminal relay model captures the trade-off between repeater-assisted and repeaterless communication strategies. We propose a decode-forward coding scheme and analyze both entanglement-assisted and unassisted scenarios. Our framework allows for different entanglement topologies between the transmitter, the relay and the destination receiver, recovering known results on entanglement-assisted and unassisted communication. Furthermore, we discuss the interpretation of coding with quantum side information. These findings serve as a stepping stone for the design of secure, efficient, and reliable quantum networks and the practical realization of quantum repeaters and long-range quantum key distribution.
@@ I Introduction
From a security standpoint, a the US National Security Agency (NSA) has highlighted several challenges regarding the current applicability of quantum cryptography [13]. Among these challenges is the need for trusted intermediate stations in current implementations of quantum key distribution (QKD). As pointed out by Renner and Wolf  [14], this will change once quantum repeaters are employed. These devices work entirely on the quantum level and hence do not need to be trusted. Therefore, using quantum repeaters will prevent security risks from insider threats. Understanding the relaying of quantum information over quantum channels can serve as a significant milestone on the road to developing quantum repeaters.
@@
In multihop networks, multiple rounds of communication take place, making synchronization a crucial requirement [25, 26]. Consider a central clock that ticks $n$ times. Between clock ticks $i-1$ and $i$, the sender and relay transmit their respective channel inputs, $A_{i}$ and $D_{i}$ (see Fig. 2). At tick $i$, the relay and destination receive the channel outputs, $E_{i}$ and $B_{i}$, respectively. A short delay is thus necessary to ensure causal operation. The classical relay channel model was originally introduced by van der Meulen [27], and later adopted as a foundation for multihop network models [28].
Causality constraints in network settings often require the use of block Markov coding [29], where the transmitter sends a sequence of blocks, and each block encodes information related to both current and previous transmissions. Quantum versions of block Markov coding have been studied in the context of classical-quantum (c-q) relay channels [30, 31], quantum communication with parameter estimation [32], and quantum cribbing scenarios [33].
An information-theoretic framework for primitive quantum relay channels was first introduced in [34], and later extended in [35] to address classical communication over fully quantum relay channels using various coding strategies. Other quantum repeater models can be found in [36, 37, 38, 39, 40].
@@
In general, the quantum communication literature can broadly be categorized into two approaches: repeater-aided and repeaterless transmission. The Pirandola–Laurenza–Ottaviani–Banchi (PLOB) bound [41] sets the fundamental rate-distance limit for repeaterless communication, i.e., without a quantum repeater. The repeater-aided transmission enables the distribution of quantum correlations across long distances, by dividing the channel into shorter segments and placing where intermediate relay terminals. Through entanglement swapping, the quantum repeaters generate entanglement and maintain coherence without violating the no-cloning theorem [42, 43].
Experimental realizations for repeaterless and repeater-aided communication were demonstrated in Refs.  [44] and Refs.  [45], respectively.
@@
In this work, we focus on the transmission of quantum information over quantum relay channels, aiming to characterize the trade-offs between repeater-assisted and repeaterless communication scenarios. For the discrete-variable case, we study the three-terminal quantum relay channel illustrated in Fig. 3. This type of relay network serves as a fundamental building block in cooperative communication  [46] and can also be used to establish entanglement between network nodes  [47]. Specifically, we consider the decode-forward strategy, in which the relay fully decodes the quantum message received from Alice, re-encodes it, and forwards it to Bob. We analyze both the entanglement-assisted and unassisted cases. For the unassisted case, we leverage the fact that communication occurs over multiple rounds, allowing the protocol to reuse entanglement generated as a by-product in block $i-1$ during block $i$.
@@
Our model also permits entanglement between the inputs of Alice and the relay (see Fig. 2), which as we show leads to improved achievable transmission rates over the case of the orthogonal links (see Ref. [36] for an example of a repeater chain composed of orthogonal links). As a special case, our results recover those of Dupuis in Ref. [48], who studied the capacity of quantum channels with side information at the transmitter. From the perspective of quantum error correction, our setup can also be interpreted such that the undisturbed reference system $R$ corresponds to the quantum data register, and the ancillary system $M$, transmitted through the communication channel, enables subsequent error syndrome extraction.
@@
The paper is organized as follows. In Sec. II, we present preliminary definitions and the quantum relay channel model. Sec. III describes the coding framework for strictly causal quantum relaying. Our main results for the decode-forward strategy, including assisted and unassisted scenarios, are given in Sec. IV. Sec. V provides an overview of the fully quantum Slepian-Wolf (FQSW) protocol, a central tool in our analysis. Detailed proofs of the achievable rates appear in Sec. VI and Sec. VII. Finally, Sec. VIII provides a summary and concluding remarks.
@@ II Preliminaries and Notation
The outcome distribution of a quantum measurement can be described using a positive operator-valued measure (POVM), which is a collection of positive semi-definite operators $\{\Delta_{j}\}$ such that $\sum_{j}\Delta_{j}=\mathbbm{1}$, where $\mathbbm{1}$ is the identity operator. According to the Born rule, if the system is in the state $\rho$, the probability of observing outcome $j$ is given by $\Pr(j)=\operatorname{Tr}(\Delta_{j}\rho).$. The quantum entropy of a density operator $\rho$ is defined as $H(\rho)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}-\operatorname{Tr}(\rho\log\rho),$. which reduces to the Shannon entropy of the eigenvalue distribution of $\rho$. For a quantum state $\rho_{AB}$ defined over the joint system $A$ and $B$ with $\rho_{AB}\in\mathcal{D}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$, the quantum mutual information is defined as
@@
where $\rho_{A}=\operatorname{Tr}_{B}(\rho_{AB})$ and $\rho_{B}=\operatorname{Tr}_{A}(\rho_{AB})$ are the reduced density operators. The conditional quantum entropy is given by $H(A|B)_{\rho}=H(\rho_{AB})-H(\rho_{B})$, and the coherent information is defined as $I(A\rangle B)_{\rho}=H(\rho_{B})-H(\rho_{AB})_{\rho}=-H(A|B)_{\rho}$.
We use the following notation for state approximation: We write $\rho\mathrel{\stackrel{{\scriptstyle\mathclap{\textstyle\varepsilon}}}{{\approx}}}\sigma$ iff ${\|\rho-\sigma\|}\leq\varepsilon$. Uhlmann’s theorem is a useful tool in this work: If the reduced states of ${|{\varphi}\rangle}_{AB}$ and ${|{\psi}\rangle}_{AC}$ satisfy $\varphi_{A}\mathrel{\stackrel{{\scriptstyle\mathclap{\textstyle\varepsilon}}}{{\approx}}}\psi_{A}$, then there exists a partial isometry $V_{C\rightarrow B}$ such that
${|{\varphi}\rangle}_{AB}\mathrel{\stackrel{{\scriptstyle\mathclap{\textstyle 2\sqrt{\varepsilon}}}}{{\approx}}}V_{C\rightarrow B}{|{\psi}\rangle}_{AC}.$
In this work, we consider a fully quantum relay channel as a model for the three-terminal network depicted in Fig. 3. A quantum relay channel is described by a completely positive trace-preserving (CPTP) map $\mathcal{N}_{AD\rightarrow BE}$, where $A$, $D$, $B$, and $E$ correspond to the sender’s transmitter, the relay’s transmitter, the destination receiver, and the relay’s receiver, respectively (see Fig. 3). We assume that the channel is memoryless. Specifically, if the sender and relay inputs are given by the systems $A^{n}=(A_{1},A_{2},\ldots,A_{n})$ and $D^{n}=(D_{1},D_{2},\ldots,D_{n})$ across $n$ channel uses, then the joint input state $\rho_{A^{n}D^{n}}$ evolves under the tensor product channel $\mathcal{N}_{A^{n}D^{n}\rightarrow B^{n}E^{n}}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\left(\mathcal{N}_{AD\rightarrow BE}\right)^{\otimes n}$. The specific communication setting will be formalized in Sec. III, where we assume that the relay encodes information in a strictly causal manner. That is, at each time step $i$, the relay first transmits $D_{i}$ and only then receives $E_{i}$.
@@ III Coding Scheme
Consider a quantum relay channel $\mathcal{N}_{AD\rightarrow BE}$. We define a code for the transmission of quantum information via $\mathcal{N}_{AD\rightarrow BE}$, from Alice to Bob. A $(2^{nQ},n)$ quantum code for $\mathcal{N}_{AD\rightarrow BE}$ is defined by:
  • Finite Hilbert space $\mathcal{H}_{M}$ of dimension $\dim(\mathcal{H}_{M})=2^{nQ}$.
@@
The communication scheme is illustrated in Fig. 2. The systems $A^{n}$ and $D^{n}$ represent the transmissions from the sender and the relay, respectively, while $E^{n}$ and $B^{n}$ denote the received outputs at the relay and the receiver. The system $R$ represents the complementary system, to which the state in the relay communication scheme may be entangled. The joint state at the beginning of the protocol is described by the pure state ${|{\Psi}\rangle}\in\mathcal{D}(\mathcal{H}_{M}\otimes\mathcal{H}_{R})$. The system Alice, representing the sender, selects a quantum system $\rho_{M}$ which is intended for Bob, representing the destination receiver. The encoding is done by Alice as $\rho_{M}\rightarrow\mathcal{E}_{M\rightarrow A^{n}}(\rho_{M})$, followed by the transmission of the $A^{n}$ systems over $n$ channel uses.
@@
The first time step $i=1$ requires an "initial condition" for the encoding map $\Gamma^{(1)}$ as:
@@
In what follows, we employ the notation $A^{n}=A^{q-1}_{1}\bigcup A^{n}_{q}$, where $A^{i}_{j}=\{A_{p}\}^{i}_{p=j}$. For example, taking $q=2$, we split $A^{n}$ into $A_{1}$ and $A_{2}^{n}\equiv\{A_{p}\}_{p=2}^{n}$. At $i=1$, the relay receives $E_{1}$ with the state, after the action of the quantum channel:
$\rho_{RB_{1}E_{1}A_{2}^{n}}=\left[\mathrm{id}_{R}\otimes\mathcal{N}_{AD\rightarrow BE}\otimes\mathrm{id}_{A_{2}^{n}}\right]\rho_{RA_{1}D_{1}A_{2}^{n}}.$ (2)
@@
$\rho_{RB_{2}E_{2}A_{3}^{n}B_{1}\bar{E}_{1}}=\\ \left[\mathrm{id}_{R}\otimes\mathcal{N}_{AD\rightarrow BE}\otimes\mathrm{id}_{A_{3}^{n}B_{1}\bar{E}_{1}}\right]\rho_{RA_{2}D_{2}A_{3}^{n}B_{1}\bar{E}_{1}}.$ (4)
@@
$\rho_{R\bar{E}^{n-1}D_{n}A_{n}^{n}B^{n-1}}=\\ \left[\mathrm{id}_{R}\otimes\Gamma_{E_{n-1}\bar{E}^{n-2}\rightarrow\bar{E}^{n-1}D_{n}}^{(n)}\otimes\mathrm{id}_{A_{n}B^{n-1}}\right]\\ (\rho_{RE_{n-1}\bar{E}^{n-2}A_{n}B^{n-1}}),$ (5)
@@
The output state for Bob (destination receiver) is the reduced state $\rho_{B^{n}}=\operatorname{Tr}_{RE_{n}\bar{E}^{n-1}}\left[\rho_{RB^{n}E_{n}\bar{E}^{n-1}}\right]$. Bob receives the channel output state $\rho_{B^{n}}$ over the output system $B^{n}$, applies the decoding channel $\mathcal{P}_{B^{n}\to\hat{M}}$ and recovers $\hat{M}$ as the estimation of the quantum system sent by Alice. The error of the code is:
$e^{(n)}(\Psi,\mathscr{C})=\frac{1}{2}||\Psi_{RM}-\Psi_{R\hat{M}}||_{1}.$ (7)

### v2

@@ Abstract
Quantum relays are central to both quantum communication and distributed quantum computing, enabling long-distance transmission and modular architectures. Unlike classical repeaters, quantum repeaters preserve coherence without amplifying quantum information, relying on entanglement swapping and quantum error correction to overcome loss and decoherence. In this work, we investigate the transmission of quantum information via quantum relay channels. Our three-terminal relay model captures the trade-off between repeater-assisted and repeaterless communication strategies. Specifically, we propose a partial decode-forward strategy, in which quantum “message system" consists of two components. The first component is decoded by the relay and then sent to the destination receiver, whereas the second component is decoded by the destination receiver without the relay’s help. We analyze both entanglement-assisted and unassisted scenarios. As a special case, the full decode-forward strategy is recovered, with the relay decoding, re-encoding, and forwarding the entire message. Our framework allows for different entanglement topologies between the transmitter, the relay and the destination receiver, recovering known results on entanglement-assisted and unassisted communication. Furthermore, we discuss the interpretation of coding with quantum side information. These findings provide a foundation for designing secure, efficient, and reliable quantum networks and for realizing practical quantum repeaters and long-range quantum key distribution.
@@ I Introduction
Part of this work has been accepted and will be presented at the 61st Allerton Conference on Communication, Control, and Computing, Urbana, IL, September 2025.
@@
From a security standpoint, a the US National Security Agency (NSA) has highlighted several challenges regarding the current applicability of quantum cryptography [13]. Among these challenges is the need for trusted intermediate stations in current implementations of quantum key distribution (QKD). As pointed out by Renner and Wolf  [14, 15], this will change once quantum repeaters are employed [16]. These devices work entirely on the quantum level and hence do not need to be trusted. Therefore, using quantum repeaters will prevent security risks from insider threats. Understanding the relaying of quantum information over quantum channels can serve as a significant milestone on the road to developing quantum repeaters.
@@
In general, the quantum communication literature can broadly be categorized into two approaches: repeater-aided and repeaterless transmission. The Pirandola–Laurenza–Ottaviani–Banchi (PLOB) bound [27] sets the fundamental rate-distance limit for repeaterless communication, i.e., without a quantum repeater. The repeater-aided transmission enables the distribution of quantum correlations across long distances, by dividing the channel into shorter segments and placing where intermediate relay terminals. Through entanglement swapping, the quantum repeaters generate entanglement and maintain coherence without violating the no-cloning theorem [28, 29]. Experimental realizations for repeaterless and repeater-aided communication were demonstrated in Refs.  [30] and  [31], respectively.
In multihop networks, multiple rounds of communication take place, making synchronization a crucial requirement [32, 33]. Consider a central clock that ticks $n$ times. Between clock ticks i1i-1 and $i$, the sender and relay transmit their respective channel inputs, $A_{i}$ and $D_{i}$ (see Fig. 2). At tick $i$, the relay and destination receive the channel outputs, $E_{i}$ and $B_{i}$, respectively. A short delay is thus necessary to ensure causal operation. The classical relay channel model was originally introduced by van der Meulen [34], and later adopted as a foundation for multihop network models [35].
Causality constraints in network settings often require the use of block Markov coding [36], where the transmitter sends a sequence of blocks, and each block encodes information related to both current and previous transmissions. Quantum versions of block Markov coding have been studied in the context of classical-quantum (c-q) relay channels [37, 38], quantum communication with parameter estimation [39], and quantum cribbing scenarios [40]. An information-theoretic framework for primitive quantum relay channels was first introduced in [41], and later extended in [42] to address classical communication over fully quantum relay channels using various coding strategies. Other quantum repeater models can be found in [43, 44, 45, 46, 47].
@@
In this work, we focus on the transmission of quantum information over quantum relay channels, aiming to characterize the trade-offs between repeater-assisted and repeaterless communication scenarios. For the discrete-variable case, we study the three-terminal quantum relay channel illustrated in Fig. 3. This type of relay network serves as a fundamental building block in cooperative communication  [48] and can also be used to establish entanglement between network nodes  [49].
Specifically, we consider the partial decode-forward strategy, in which quantum “message system" consists of two components. The first component is decoded by the relay and then sent to the destination receiver, whereas the second message component is decoded by the destination receiver without the relay’s help. We derive achievable rates, both with and without entanglement assistance. As a special case, we obtain an achievable rate corresponding to a full decode-forward strategy, in which the relay decodes the entire information, re-encodes it, and forwards it to the destination receiver. In the unassisted case, we leverage the fact that communication occurs over multiple rounds, allowing the protocol to reuse entanglement generated as a by-product in block i1i-1 during block $i$.
Explicitly, for the unassisted case of the partial decode-forward strategy we prove that the quantum capacity of the quantum relay channel $\mathcal{N}_{AD\rightarrow BE}$, satisfies
We show that the quantum rate $Q$ below is achievable for the quantum relay channel:
$Q=\max_{\sigma}\bigl{[}\min\bigl{\{}I(A_{1}A_{2}\rangle B)_{\omega}\,,\\ \;I(A_{1}\rangle E)_{\omega}+I(A_{2}\rangle BA_{1})_{\omega}\bigr{\}}\bigr{]}$ (1)
where $I(A\rangle B)_{\rho}$ denotes the coherent information corresponding to a quantum state $\rho_{AB}$. Here, $A$, $D$, $E$, and $B$ represent the sender’s transmitter, the relay’s transmitter, the relay’s receiver, and the destination receiver, respectively (see Fig. 3). The maximum is taken over all $|{\sigma_{A_{1}A_{2}AD}}\rangle$}, where $A_{1}$ and $A_{2}$ are ancillas that represent the quantum message components mentioned above. Intuitively, $A_{1}$ represents Alice’s share of entanglement with the relay, while $A_{2}$ is Alice’s share of entanglement with Bob. Due to entanglement monogamy [50], there is a tradeoff between the entanglement that Alice generates with the relay and with the destination decoder.
@@
Our model also permits entanglement between the inputs of Alice and the relay (see Fig. 2), which as we show leads to improved achievable transmission rates over the case of the orthogonal links (see Ref. [43] for an example of a repeater chain composed of orthogonal links). The capacity bound with entanglement assistance resembles that of Dupuis Ref. [51], who studied the entanglement-assisted capacity of quantum channels with side information at the transmitter.
While QEC has advanced significantly [52, 53, 54], practical bottlenecks remain, especially the need for rapid, accurate decoding to avoid syndrome backlogs that can exponentially slow computation [55, 56]. A promising approach is to parallelize syndrome extraction across distributed devices, where ancillary registers are connected through quantum relays enabling reliable state transfer via entanglement swapping and teleportation [57, 58, 59]. In that context, our setup can also be interpreted as follows. Consider the communication setup depicted in Fig. 2, the undisturbed reference system $R$ corresponds to the quantum data register, and the ancillary system $M$, transmitted through the repeater-assisted communication channel enables subsequent error syndrome extraction. This distributed approach could enhance the performance of fault-tolerant architectures, paving the way toward more scalable, higher-fidelity quantum error correction in large quantum systems [60, 61].
@@
The paper is organized as follows. In Sec. II, we present preliminary definitions and the quantum relay channel model. Sec. III describes the coding framework for strictly causal quantum relaying. Our main results for the partial and full decode-forward strategies, including assisted and unassisted scenarios, are given in Sec. IV. Then, in Sec. V we give an intuitive analysis of our main results. Sec. VI provides an example based on the quantum erasure channel. Sec. VII provides a summary and concluding remarks. Finally, Appendix A overviews the fully quantum Slepian-Wolf (FQSW) protocol, central to our analysis, while detailed proofs of the achievable rates are given in Appendix B and Appendix C.
@@ II Preliminaries and Notation
The outcome distribution of a quantum measurement can be described using a positive operator-valued measure (POVM), which is a collection of positive semi-definite operators $\{\Delta_{j}\}$ such that $\sum_{j}\Delta_{j}=\mathbbm{1}$, where $\mathbbm{1}$ is the identity operator. According to the Born rule, if the system is in the state $\rho$, the probability of observing outcome $j$ is given by $\Pr(j)=\operatorname{Tr}(\Delta_{j}\rho).$ The quantum entropy of a density operator $\rho$ is defined as $H(\rho)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}-\operatorname{Tr}(\rho\log\rho),$ which reduces to the Shannon entropy corresponding to the eigenvalue distribution of $\rho$. For a bipartite quantum state $\rho_{AB}\in\mathcal{D}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$, the quantum mutual information is defined as
@@
where $\rho_{A}=\operatorname{Tr}_{B}(\rho_{AB})$ and $\rho_{B}=\operatorname{Tr}_{A}(\rho_{AB})$ are the reduced density operators. The conditional quantum entropy is defined as $H(A|B)_{\rho}=H(\rho_{AB})-H(\rho_{B})$, and the coherent information as $I(A\rangle B)_{\rho}=H(\rho_{B})-H(\rho_{AB})_{\rho}=-H(A|B)_{\rho}$.
We use the following notation for state approximation: We write $\rho\mathrel{\stackrel{{\scriptstyle\mathclap{\textstyle\varepsilon}}}{{\approx}}}\sigma$ iff ${\|\rho-\sigma\|}_{1}\leq\varepsilon$. Uhlmann’s theorem is a useful tool in this work: consider the pure states ${|{\varphi}\rangle}_{AB}$ and ${|{\psi}\rangle}_{AC}$, where ${|{\psi}\rangle}_{AC}$ is non-normalized. If the reduced states satisfy $\varphi_{A}\mathrel{\stackrel{{\scriptstyle\mathclap{\textstyle\varepsilon}}}{{\approx}}}\psi_{A}$, then there exists a partial isometry $V_{C\rightarrow B}$ such that
${|{\varphi}\rangle}_{AB}\mathrel{\stackrel{{\scriptstyle\mathclap{\textstyle 4\sqrt{\varepsilon}}}}{{\approx}}}V_{C\rightarrow B}{|{\psi}\rangle}_{AC}.$
Note that this version of the Uhlmann’s theorem is essentially a generalization of the original theorem to the case where the purifications are allowed to be non-normalized, introducing an additional factor of $2$ into the above approximation (see Lemma 5 in Ref. [62]).
In this work, we consider a fully quantum relay channel as a model for the three-terminal network depicted in Fig. 3. A quantum relay channel is described by a completely-positive trace-preserving (CPTP) map $\mathcal{N}_{AD\rightarrow BE}$, where $A$, $D$, $B$, and $E$ correspond to the sender’s transmitter, the relay’s transmitter, the destination receiver, and the relay’s receiver, respectively (see Fig. 3). We assume that the channel is memoryless. Specifically, if the sender and relay inputs are given by the systems $A^{n}=(A_{1},A_{2},\ldots,A_{n})$ and $D^{n}=(D_{1},D_{2},\ldots,D_{n})$ across $n$ channel uses, then the joint input state $\rho_{A^{n}D^{n}}$ evolves under the tensor product channel $\mathcal{N}_{A^{n}D^{n}\rightarrow B^{n}E^{n}}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\left(\mathcal{N}_{AD\rightarrow BE}\right)^{\otimes n}$. The specific communication setting will be formalized in Sec. III, where we require that the relay encodes information in a strictly causal manner. That is, at each time step $i$, the relay first transmits $D_{i}$ and only then receives $E_{i}$.
@@ III Coding Scheme
Consider a quantum relay channel $\mathcal{N}_{AD\rightarrow BE}$. We define a code for entanglement generation via $\mathcal{N}_{AD\rightarrow BE}$, from Alice to Bob. A $(2^{nQ},n)$ quantum code for $\mathcal{N}_{AD\rightarrow BE}$ is defined by:
  • A Hilbert space $\mathcal{H}_{M}$ of dimension $\dim(\mathcal{H}_{M})=2^{nQ}$.
@@
The communication scheme is illustrated in Fig. 2. The systems $A^{n}$ and $D^{n}$ represent the transmissions from the sender and the relay, while $E^{n}$ and $B^{n}$ denote the received outputs at the relay and the receiver, respectively,. At the beginning of the protocol, Alice holds a “message system" $M$, in a state $\rho_{M}$. Let ${|{\Psi}\rangle}\in\mathcal{D}(\mathcal{H}_{M}\otimes\mathcal{H}_{R})$ be a purification. The system $R$ represents the complementary system, to which the state in the relay communication scheme may be entangled. In the quantum communication protocol, Alice would like to send $M$ to Bob. The encoding is done by Alice by a mapping of $\rho_{M}\mapsto\mathcal{E}_{M\to A^{n}}(\rho_{M})$, followed by the transmission of the $A^{n}$ systems over $n$ channel uses.
@@
The first time step $i=1$ requires an “initial condition" for the encoding map $\Gamma^{(1)}$ as:
@@
In what follows, we employ the notation $A_{i}^{j}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}A_{i}A_{i+1}\ldots A_{j}$ for $i<j$. At $i=1$, the relay receives $E_{1}$ with the state, after the action of the quantum channel:
$\rho_{RB_{1}E_{1}A_{2}^{n}}=\left[\mathrm{id}_{R}\otimes\mathcal{N}_{AD\rightarrow BE}\otimes\mathrm{id}_{A_{2}^{n}}\right]\rho_{RA_{1}D_{1}A_{2}^{n}},$ (3)
where $A^{n}=A_{1}A_{2}^{n}$.
@@
$\rho_{RB_{2}E_{2}A_{3}^{n}B_{1}\bar{E}_{1}}=\\ \left[\mathrm{id}_{R}\otimes\mathcal{N}_{AD\rightarrow BE}\otimes\mathrm{id}_{A_{3}^{n}B_{1}\bar{E}_{1}}\right]\rho_{RA_{2}D_{2}A_{3}^{n}B_{1}\bar{E}_{1}},$ (5)
and so on.
@@
$\rho_{R\bar{E}^{n-1}D_{n}A_{n}^{n}B^{n-1}}=\\ \left[\mathrm{id}_{R}\otimes\Gamma_{E_{n-1}\bar{E}^{n-2}\rightarrow\bar{E}^{n-1}D_{n}}^{(n)}\otimes\mathrm{id}_{A_{n}B^{n-1}}\right]\cdot\\ (\rho_{RE_{n-1}\bar{E}^{n-2}A_{n}B^{n-1}}),$ (6)
@@
The output state for Bob (destination receiver) is the reduced state $\rho_{B^{n}}=\operatorname{Tr}_{RE_{n}\bar{E}^{n-1}}\left[\rho_{RB^{n}E_{n}\bar{E}^{n-1}}\right]$. Bob receives the channel output state $\rho_{B^{n}}$ of the output system $B^{n}$, applies the decoding channel $\mathcal{P}_{B^{n}\to\hat{M}}$ and recovers $\hat{M}$ as the estimation of the quantum system sent by Alice. This results in the final state $\rho_{R\hat{M}}=\mathcal{P}_{B^{n}\to\hat{M}}(\rho_{RB^{n}})$. The error of the code is:
[…]
A quantum transmission rate $Q>0$ is called achievable if for every $\varepsilon,\delta>0$ and sufficiently large $n$, there exists a $(2^{n(Q-\delta)},n)$ code such that $e^{(n)}(\Psi,\mathscr{C})\leq\varepsilon$ for all ${|{\Psi_{RM}}\rangle}\in\mathcal{H}_{M}\otimes\mathcal{H}_{R}$. The quantum capacity $C_{\text{Q}}(\mathcal{N})$ is defined as the supremum of achievable quantum rates.
One may focus on the task of entanglement generation. In this case, an entanglement-generation rate $Q_{\text{EG}}>0$ is called achievable if for every $\varepsilon,\delta>0$ and sufficiently large $n$, there exists a $(2^{n(Q_{\text{EG}}-\delta)},n)$ code such that $e^{(n)}(\Phi,\mathscr{C})\leq\varepsilon$, where $|{\Phi_{RM}}\rangle$} is the canonical maximally entangled state on $\mathcal{H}^{\otimes 2}_{M}$. The entanglement-generation capacity $C_{\text{EG}}(\mathcal{N})$ is defined as the supremum of achievable entanglement-generation rates.